You are evaluating the purchase of a vehicle for your business. You've decided that the best choice is a car that will cost you $35,000, but you're uncertain how long you should plan on holding the car before you replace it. The table below lists the running costs and salvage value of the vehicle for each year. 1 2 3 4 5 Running Costs -3000 -3500 -4000 -4500 -5000 Salvage Value 25000 20000 15000 10000 5000 What is the annual equivalent cost of replacing the vehicle every 3 years? Assume your cost of capital is 16.9%. Enter your answer to the nearest cent. Ignore taxes. (Your answer will be a cost, and therefore a negative number. Don't forget the minus sign.)
Answer #1

First, we calculate the NPV

Year 0 1 2 3
Initial Cost -35000
Running Cost -3000 -3500 -4000
Salvage value 15000
Net Cost -35000 -3000 -3500 11000
NPV at 16.9% $ -33,241.75

Equivalent annual cost = NPV * r/ (1-(1+r)^-n ) = -33,241.75*0.169/(1-1.169^-3) = -15,019.97

Equivalent annual cost = -$15,019.97 (negative indicates cost)